# Convergence of Pivoted Cholesky for Lipschitz Kernels

### v1

@@ 1. Introduction
The pivoted Cholesky algorithm is best viewed as an iterative low-rank approximation scheme, rather than as a factorization. Formally, let $\Omega\subset\mathbb{R}^{d}$ be a compact domain. Given a SPD kernel [^1] A kernel is called SPD if for any finite set of points $x_{1},\ldots,x_{n}\in\Omega$, the matrix $\big{[}K(x_{j},x_{k})\big{]}_{j,k=1}^{n}$ is SPD. $K:\Omega\times\Omega\rightarrow\mathbb{R}$, the first step of the pivoted Cholesky algorithm sets $R_{0}=K$ and selects a point $z_{1}$ (which we call a pivot) based on prescribed selection criteria, which we call the pivoting strategy. Then, the algorithm constructs a rank-$1$ approximation of $K$ as
@@
which interpolates $K$ whenever $x=z_{1}$ or $y=z_{1}$ and the first residual $R_{1}=K-K_{1}$ is obtained. We iteratively apply this procedure. At the $n\geq 1$ step, we select a pivot point $z_{n}$ according to the pivot strategy and compute the rank-$n$ approximant of $K$ by adding a rank-$1$ kernel to a previous approximant, i.e.,
@@ 1.1. Our contributions
There are also matrix counterparts to Theorem 3.2, which we find intriguing, giving us a decay rate for the Cholesky decomposition applied to structured matrices. Let $A$ be an $m\times m$ SPD matrix. Then, for $1<n\leq m$, we have (see Corollary 6.1)
@@ 4.3. Discrete pivoting
Thus, choosing a grid with spacing $\eta$ such that $4L\eta\leq(1-\delta)\|R_{i-1}\|_{\infty}$ guarantees that the grid-based pivot achieves at most a $\delta$ pivoting mistake at step $i$. In practice, a simple strategy is to start with a coarse grid and refine it adaptively so that $\delta$ remains near $1$ while keeping evaluations inexpensive.
@@ 6.1. Matrix counterparts of the kernel results
This quantity is the matrix analogue of the Lipschitz constant $L$ from the kernel setting: it measures how rapidly the entries of $A$ vary along each row as the column index changes. Applying the Cholesky algorithm with complete pivoting to $A$ then yields the following discrete counterpart of Theorem 3.2.
@@
Let $A$ be an $m\times m$ SPD matrix and define $G_{A}$ as in eq. 10. Let $A_{n}$ be the rank-$n$ approximation obtained by the Cholesky decomposition with complete pivoting. Then, for $1<n\geq m$, we have

### v2

@@ 1. Introduction
The pivoted Cholesky algorithm is best viewed as an iterative low-rank approximation scheme, rather than as a factorization. Formally, let $\Omega\subset\mathbb{R}^{d}$ be a compact domain. Given a SPD kernel [^1] A kernel is called SPD if for any finite set of points $x_{1},\ldots,x_{n}\in\Omega$, the matrix $\big[K(x_{j},x_{k})\big]_{j,k=1}^{n}$ is SPD. $K:\Omega\times\Omega\rightarrow\mathbb{R}$, the first step of the pivoted Cholesky algorithm sets $R_{0}=K$ and selects a point $z_{1}$ (which we call a pivot) based on prescribed selection criteria, which we call the pivoting strategy. Then, the algorithm constructs a rank-11 approximation of $K$ as
@@
which interpolates $K$ whenever $x=z_{1}$ or $y=z_{1}$ and the first residual $R_{1}=K-K_{1}$ is obtained. We iteratively apply this procedure. At the $n\geq 1$ step, we select a pivot point $z_{n}$ according to the pivot strategy and compute the rank-$n$ approximant of $K$ by adding a rank-11 kernel to a previous approximant, i.e.,
@@
There are also matrix counterparts to Theorem 3.2, which we find intriguing, giving us a decay rate for the Cholesky decomposition applied to structured matrices. Let AA be an $m\times m$ SPD matrix. Then, for $1<n\leq m$, we have (see Corollary 6.1)
@@ 4.3. Discrete pivoting
Thus, choosing a grid with spacing $\eta$ such that $4L\eta\leq(1-\delta)\|R_{i-1}\|_{\infty}$ guarantees that the grid-based pivot achieves at most a $\delta$ pivoting mistake at step ii. In practice, a simple strategy is to start with a coarse grid and refine it adaptively so that $\delta$ remains near 11 while keeping evaluations inexpensive.
@@ 6.1. Matrix counterparts of the kernel results
This quantity is the matrix analogue of the Lipschitz constant $L$ from the kernel setting: it measures how rapidly the entries of AA vary along each row as the column index changes. Applying the Cholesky algorithm with complete pivoting to AA then yields the following discrete counterpart of Theorem 3.2.
@@
Let AA be an $m\times m$ SPD matrix and define $G_{A}$ as in eq. 10. Let $A_{n}$ be the rank-$n$ approximation obtained by the Cholesky decomposition with complete pivoting. Then, for $1<n\geq m$, we have